# Training Robust Graph Neural Networks by Modeling Noise Dependencies

### v1

@@ Appendix F Additional Experimental Results
F.1 Robustness evaluation under existing noise scenarios
In Fig 11, 11, and 11, we show Da-Gnn exhibits comparable or better performance than baselines under other noise scenarios, commonly considered in this research field.
F.2 Comparison with the naive combination of existing works
@@
F.3 Complexity Analysis
Large Scale Graph. To demonstrate the scalability of Da-Gnn, we consider a larger graph dataset, ogbn-arxiv (Hu et al., 2020). Table 1 clearly illustrates that Da-Gnn exhibits superior scalability and robustness in comparison to other baseline methods.
Table 7: Training time comparison on Cora dataset under DANG 50%.
Total (sec) 20.9 702.1 159.9 53.3 0.8 100.3 118.7 86.3 46.3
per epoch (sec) 0.04 1.77 0.16 - 0.004 0.20 0.18 0.11 0.09
F.4 Sensitivity Analysis
In this section, we analyze the sensitivity of the coefficient $\lambda_{1}$ and $\lambda_{2}$ in Eqn 3, $\theta$, and $\gamma$. To be specific, we increase $\lambda_{1}$ value from $\{0.0,0.003,0.03,0.3,3\}$, $\lambda_{2}$ value from $\{0.0,0.003,0.03,0.3\}$, $\theta$ from $\{0.1,0.2,0.3\}$, and $\gamma$ from $\{0,1\}$. We then evaluate the node classification accuracy of Da-Gnn under DANG.
  • In Fig 12(a), we notice that Da-Gnn consistently surpasses the state-of-the-art baseline, EvenNet, regardless of the $\lambda_{1}$ value, demonstrating the robustness of Da-Gnn. Furthermore, we observe that the performance significantly drops when $\lambda_{1}=0$. This highlights the importance of modeling the causal relationship $A\leftarrow(X,\epsilon,Z_{A})$ for robustness under DANG, as $\lambda_{1}$ is directly related to the loss term $\mathcal{L}_{\text{edge-rec}}$, i.e., $-\mathbb{E}_{Z_{A}}\mathbb{E}_{\epsilon}\left[\log(p_{\theta_{1}}(A|X,\epsilon,Z_{A}))\right]$.
  • In Fig 12(b), we observe that Da-Gnn generally outperforms the sota baseline regardless of the value of $\lambda_{2}$, indicating the stability of Da-Gnn. Moreover, we can see a performance decrease when $\lambda_{2}=0$. This observation suggests that the regularization on the inferred latent node label $Z_{Y}$ using the inferred latent structure $Z_{A}$ effectively handles the noisy labels. This conclusion is drawn from the fact that $\lambda_{2}$ is directly linked to the loss term $\mathcal{L}_{\text{hom}}$, i.e., $kl(q_{\phi_{3}}(Z_{Y}|X,A)||p(Z_{Y}))$.
  • In Fig 12(c), we observe that Da-Gnn consistently outperforms the state-of-the-art baseline, EvenNet, across all values of $\theta$, demonstrating the robustness of the prediction regularization method in Eqn2.
  • In Fig 12(d), we observe that Da-Gnn consistently surpasses the state-of-the-art baseline, EvenNet, across all values of $\gamma$, emphasizing the stability of regularizing the inferred $Z_{A}$ in modeling $q_{\phi_{1}}(Z_{A}|X,A)$.
@@ F.4 Sensitivity Analysis
3:  $\mathcal{V}^{\text{noisy}}\leftarrow$ Randomly sample a $\eta\%$ subset of nodes
4:  $\mathbf{X}^{\text{noisy}}\leftarrow\mathbf{X}$

### v2

@@ F.3.2 Variants of Real-world DANG
F.4 Qualitative Analysis
Furthermore, to verify the distinction between the noisy and clean scenarios, We conduct a non-parametric analysis, Mann–Whitney U test, which require no distributional assumptions. The results are as follows:
  • Fig 13(a): Statistic=$62337852.0$, p-value=$0.0$
  • Fig 13(b): Statistic=$40277922.0$, p-value=$0.0$.
Note that we found the scipy.stats package displays p-values as zero when they are extremely low. Therefore, we reported the corresponding test statistics with p-values. The results indicate highly significant differences between the groups.
F.5 Comparison with the Naive Combination of Existing Works
@@
[…]